OpenStudy (anonymous):

what is the square root of 20?

OpenStudy (anonymous):

\[2*\sqrt{5}\]

OpenStudy (anonymous):

20 = 4*5 2 sqrt 5

OpenStudy (hoblos):

√20 = √(4*5) = 2√5

OpenStudy (anonymous):

sqrt4 X sqrt5 which equals 2 sqrt(5)

OpenStudy (anonymous):

\[2\sqrt{5}\] or =4.4721